OpenStudy (anonymous):

Find all the zeroes of each function f(x) = x3 - 4x2 + 6x - 4

OpenStudy (anonymous):

\[f(x)=x^3-4x^2+6x-4\]

OpenStudy (anonymous):

maybe you can factor

OpenStudy (anonymous):

my online textbook is down so i cant study anything i juss have to take quizes and hope for the best so i really need help idk even kno how to start this problem

OpenStudy (anonymous):

(x-2)*(x^2-2*x+2) well it factors to this and you can use the quadriatic formula for the the x^2 part. from the first group of parenthesis you can tell 2 is a zero :)

OpenStudy (anonymous):

i honestly don't understand anything about this how is 2 a zero?

OpenStudy (anonymous):

you set your equation to 0 :)

OpenStudy (anonymous):

so i write the equation out and then put = 0 at the end

OpenStudy (anonymous):

\[{x}^{3}-4\,{x}^{2}+6\,x-4=0 then solve for x \]

OpenStudy (anonymous):

ohk hold on

OpenStudy (anonymous):

how do i do 4x^2?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

you have to factor it :)

OpenStudy (anonymous):

uhm i kno wat i got is wrong sooooo yea i got -4/9

OpenStudy (anonymous):

no it is not factor your equation you'll get (x-2)*(x^2-2*x+2) so 2 is a zero and then use quadratic formula for (x^2-2*x+2)

OpenStudy (anonymous):

how do you kno 2 is 0?

OpenStudy (anonymous):

ooh jane if you allowed to use graphing calculator you can also graph it and find the zeros visually :) I know 2 is a zero because once you factor you can set each (...) inside the parentheses to = 0

OpenStudy (anonymous):

for real? how? i have a graphing calculator and i can use it

OpenStudy (anonymous):

for example \[{x}^{3}+6\,{x}^{2}+11\,x+6 =0\] is (x+1)*(x+2)*(x+3) factored out you can do this to each ( ) group (x+1)=0 (x+2)=0 (x+3)=0 x= -1, -2, and -3 :)

OpenStudy (anonymous):

ooh what kind of graphing calc do you have @janessia

OpenStudy (anonymous):

TI-83 Plus

OpenStudy (anonymous):

ooh :)

OpenStudy (anonymous):

how do i factor something out

OpenStudy (anonymous):

oh i forgot to give you my answer choices 2i,-2i,3 -4,1+2i,1-2i,,2 5,3i.-3i,-5 4,2-i,2+i 2,1+i,1-i

OpenStudy (anonymous):

well in mines when you in the graphing window you go the the f5 (math) then zeros you have to pick a low bound and higer bound :)

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

wat kind of calc do u have?

OpenStudy (anonymous):

ti 89 titanium

OpenStudy (anonymous):

oh WOWWWWWW lol

OpenStudy (anonymous):

why wow :P

OpenStudy (anonymous):

it wont work tho cuzz it doesn't give you imaginary numbers :(

OpenStudy (anonymous):

lol anyway wat do i do on the calc?

OpenStudy (anonymous):

well plug in you equation graph it when you in the window where it shows the curve find the function that lets you find the zeros it wont work tho because you are asked for imaginary roots too.

OpenStudy (anonymous):

why won't it work!!??? well wat are imaginary roots? u mean real numbers?

OpenStudy (anonymous):

non real numbers are imaginary numbers :P

OpenStudy (anonymous):

the mods have been on my butt :( i cant

OpenStudy (anonymous):

(x-2)*(x^2-2*x+2) this is factored out for you x=2 use quadratic formula to find the other 2 of (x^2-2*x+2)

OpenStudy (anonymous):

i kno the other two are ones rite?

OpenStudy (anonymous):

1+i and 1-i